# A Fast and Optimal Learning-based Path Planning Method for Planetary Rovers

### v2

@@ 4.1 Network Structure
In this section, we introduce a neural network model for path planning. The model takes as input a three-channel combination tensor, where the first channel represents a cost map as detailed in part B, with the grayscale value of each pixel indicating the cost of the robot passing through that pixel, with higher costs indicating pixels that are closer to being impassable. The second channel encodes the path planning starting point using a Gaussian representation, as described in part C. Similarly, the third channel represents the endpoint’s Gaussian encoding. The model’s output is a two-dimensional single-channel probability map, visualized as a color map in Figure. 2. Each pixel in the probability map stores a probability value ${p}$ between 0 and 1, indicating the probability that the pixel contains the optimal path. Higher probabilities imply the model is more inclined to believe the optimal path passes through that pixel.
@@
The detailed structure of the model is shown in Figure. 2. It uses an encoder-decoder architecture comprising a backbone module, pyramid pooling module and decoder module. From a unique perspective, the role of the model is equivalent to performing semantic segmentation on the map containing the start and end points, segmenting out those pixels that represent feasible paths, while we also expect the model to have a very fast inference speed. Therefore, in designing the model, we drew inspiration from excellent examples of real-time semantic segmentation models.
@@
The NNPP model employs the STDCfan2021rethinking as the backbone network. Within its multi-stage encoding architecture, the spatial resolution of feature maps is reduced by a factor of two after each STDC stage. As such, a high-level semantic description of the input map can be extracted as the final stage 5 output of the STDC backbone. The low-level features output by Stage3 and Stage4 will be fused with the high-level features in the decoder to enhance the model’s performance for path planning. Feature fusion across different semantic levels has been shown to be effective in improving performance for tasks requiring precise localization such as path planningwang2020neural , by incorporating both low-level spatial details and high-level semantic information. Therefore, in our model’s decoder, we fuse the low-level features from earlier stages with the high-level features, in order to leverage spatial details that can help produce more accurate path planning outputs.
@@
To enhance global scene contextualization, PSPNetzhao2017pyramid incorporated a pyramid pooling module (PPM) that concatenates pooling representations at multiple scales prior to the convolution layers, thereby forming both local component features and global contextual representations. Compared to the original PPM, we simplify it to use three different pyramid scales, and remove the concatenate operation between the original feature map and the upsampled features, as in the original PPM. Our simplified PPM can reduce computational cost. Specifically, we reduce the number of pyramid scales from the original 4 to only 3. Previous work has shown that features at different pyramid levels become increasingly semanticfu2019dual . Therefore, a fewer number of pyramid levels, such as 3, can capture sufficient semantic information, while reducing computational cost.
@@
Feature fusion modules are commonly employed in semantic segmentation networks to enhance feature representations. More specifically, fusing multi-level features is instrumental in achieving high segmentation accuracy. The proposed model draws inspiration from the design of pplitesegpeng2022pp , utilizing an Attention Fusion Module to amalgamate low-level features and high-level features from the decoder. The structure of the Attention Fusion Module (AFM)peng2022pp is shown in Figure. 10.
@@ 4.2 Dataset Generation
In the field of semantic segmentation for planetary surfaces, generating annotated data and training labels can be a very time-consuming taskdai2022segmarsvit . In the context of path planning tasks on lunar or Martian surfaces, obtaining the true optimal path can also be challenging. To address this issue, NNPP does not require any manually labeled data. Instead, it uses the optimal path obtained by ${A^{\star}}$ as a label to train the model. The 20-meter resolution lunar DEM dataset from CE2TMap2015xin2020geometric was used to generate a ${256\times 256}$ map. The cost of traversing each pixel was calculated using the method outlined in Eq. 1.
The classical ${A^{\star}}$ algorithm is a graph search algorithm commonly utilized for global path planningxie20202 . It is capable of finding the shortest path for a robot to travel between two predetermined positions in a known static environment. To facilitate the rover’s automatic avoidance of untraversable areas caused by hazard geomorphic terrain, this paper defines the cost function of ${A^{\star}}$ to include the traversability cost as follows:
@@
$G\left(n\right)={L_{sn}}+\omega\cdot{T_{sn}}$ (4)
Where ${{L_{sn}}=\sum_{k=1}^{n-1}{l_{k,k+1}}}$ represents the accumulated Euclidean distance that from the starting grid ${s}$ to the current grid ${n}$, and ${{T_{sn}}=\sum_{k=1}^{n-1}{T_{k,k+1}}}$ is the accumulated traversability cost between the two grids. This paper set ${H\left(n\right)}$ to be the Euclidean heuristic function.
@@ 4.3 Detailed Training Process and Evaluation Metric
Positional encoding is a powerful technique for infusing data with information regarding its position via embedding. In natural language processing (NLP), this approach is often leveraged to ensure that a model is aware of the location of words within a sentence. Inspired by this technique, NNPP makes use of positional encoding to embed location information about the starting and ending points on a map. The resulting encoded output is concatenated with the map, generating a three-channel tensor that serves as the input for the model. After encoding the starting point using sine and cosine functions to represent its absolute positionvaswani2017attention , the model’s outputs were not feasible. However, by switching to relative position encoding, more desirable results were achieved. This is because in practical path planning, the system is less concerned with absolute positions and more focused on relative distances between points. Specifically, our interest lies in the relative position of each cell in the elevation map compared to the starting point and the goal point.
@@
In this paper, a two-dimensional Gaussian distribution function centered around the starting point is used as the position encoding, with each grid in the elevation map represented by its corresponding value in the Gaussian function. This value reflects the relative position of the grid with respect to the starting point. The same approach is taken for the ending point.
Eq. 5 represents the calculation method for the Gaussian distribution function value at ${(x,y)}$. With ${\sigma}$ set to 50, ${cx}$ represents the ${x}$-coordinate of either the starting or ending point, while ${cy}$ represents the ${y}$-coordinate. Figure. 3 depicts the visualization results of Gaussian position encoding with different ${\sigma}$ values. The point coordinates are set to ${(200,200)}$. When ${\sigma}$ is too small, the effect is unsatisfactory, while when ${\sigma}$ is too large, the model also cannot extract sufficient starting point information from the encoding results. Experiments in [subsequent chapters] demonstrate that ${\sigma}$ at ${1/5}$ of the map size yields satisfactory results.
In the field of path planning, scholars have proposed many evaluation methods for path qualityliu2023path , among which average path planning time is an important metric that reflects the search efficiency of the algorithm. In elevation maps, the proposed method uses the cumulative traverse cost of a path to measure its optimality, believing that an optimal path should have the minimum cumulative traverse cost. This paper employs a learning-based approach to predict regions within a given map that contain optimal paths, whereas existing evaluation metrics are designed for a complete path rather than a region. Consequently, there is a need for designing an evaluation metric that directly assesses the precision of the model predictions. Figure. 4 illustrates the results of segmenting the model outputs according to diverse thresholds ${td}$, with the black areas representing regions assigned probabilities lower than ${td}$ by the model. The designation ${{}^{\prime}connected^{\prime}}$ denotes whether the segmented regions unite the starting point and end point. The present study decreases ${td}$ gradually from high to low values while detecting whether the two points are connected, thereby determining an optimal binary threshold and the corresponding region segmentation map. Subsequently, the total number of pixels that the model deems to contain optimal paths within the segmented regions is calculated, with the metric evaluated using the following Eq. 6. ${Z_{label}}$ represents the area of the labeled path, ${Z_{model}}$ denotes the area segmented using the adaptive threshold, and the areas are measured in number of pixels.
@@
In summary, there are three evaluation metrics: the average path planning time ${(AT)}$, the cumulative traverse cost of a path ${(CC)}$, and ${(MM)}$ as defined in Eq. 6.
@@ 5 SIMULATION EXPERIMENTS
In the present section, we delve into intricate parameter tuning and hardware environment settings, designing three experiments to demonstrate the efficacy of the proposed method. As detailed in 4.2, a training set of 100,000 samples of resolution ${256\times 256}$ was synthesized, 5000 samples were procured for elevation evaluation, and an additional 1000 samples were sequestered for performance metric validation in the present experiments. The training, elevation and validation sets were mutually exclusive from each other. The optimizer of choice for parameter optimization was AdamW with an initial learning rate of 0.001 and a weight decay of ${{10^{-4}}}$. After training for 150 epochs, the model was able to predict traversable areas between start and end points on 1000 terrain samples that it had not been exposed to previously, showcasing its generalization capabilities.
5.1 Comparative Experiments with Conventional Methods at Different Scales
The computationally intensive nature of utilizing algorithm ${A^{\star}}$ to search for optimal paths within elevation maps stems from the stipulation that ${A^{\star}}$ algorithm necessitates constructing a global search graph in advance of searching. The graph encodes the traversal costs for each pixel, incorporating the traversal costs between adjacent pixels.
@@ 5.1 Comparative Experiments with Conventional Methods at Different Scales
However, the proposed approach allows the neural network to focus on those regions that provide high value to subsequent path reconstruction algorithms such as ${RRT^{\star}}$ and ${A^{\star}}$. By reducing the spatial redundancy in the search space, the network can learn a more discriminative representation that facilitates more efficient replanning. Taking ${A^{\star}}$ as an example, constructing the search graph only within the regions output by NNPP, rather than a global search graph, as illustrated in Figure. 5, can save significant time. Moreover, as the map size increases, the advantages of the neural network approach become more pronounced. By leveraging the ability of the network to focus the search in the most promising areas, the proposed localized search graph yields increasingly superior speedup and scalability compared to exhaustive search-based algorithms as the problem size grows.
@@
The model was trained on data of size ${256\times 256}$. When evaluating the model’s performance on ${512\times 512}$ maps, as illustrated in Figure. 6, the maps were first downsampled via average pooling. Denoted by s and g are the start and goal positions respectively. Subsequently, the method described in 4.3 was applied to obtain the mask of the regions of interest to the model. Finally, these regions were resized back to the original map size via interpolation. This procedure allowed the model trained on smaller maps to generalize well to larger maps, by focusing its search only on the relevant areas while ignoring the rest. The ability to downsample the input during inference and selectively pay attention to important regions enabled by the proposed method was critical for achieving scalability. The search graph was constructed only within the black regions in Figure. 4. Subsequently, the ${A^{\star}}$ method was applied on the search graph to compute the start-to-goal path. In comparison, an exhaustive search graph was built on the original map, and the same ${A^{\star}}$ method was utilized to search for the optimal path. By constraining the search space to only the salient areas identified by the proposed method, the localized search graph approach achieved significantly higher efficiency and scalability compared to the exhaustive search baseline. The ability to prune irrelevant regions of the map based on the predicted importance masks allowed the search complexity to grow sublinearly with the map size, in contrast to the quadratic complexity of exhaustive search. This demonstrates the effectiveness of the proposed approach in focusing the search in the most promising regions to guide the path planning algorithm.
@@
As shown in the Figure. 6, the masked graph making module accepts two inputs: the model output post masking and the original map. In the original map, regions of interest as discerned by the model are retained and delineated in black while the remainder is rendered entirely white, with the white regions signifying impassable obstacles. In this manner, ${A^{\star}}$ search for the optimal path is constrained to a relatively smaller region. Provided that the specified region encompasses the optimum path, ${A^{\star}}$ will return a 100% consistent path matching the given label.
@@
As Figure. 5 illustrates, the blue bars represent the computational time incurred by the NNPP method via the workflow in Figure. 6 to derive the optimal path whereas the orange bars denote the time consumed by conventional ${A^{\star}}$ search directly on the map. Evidently, as the map size increases, the neural network more decisively shortens the search time. Figure. 5b indicates the performance metrics measured using Eq. 6 and corresponding standard deviations for the NNPP model at diverse resolutions. In larger scenes, employing the upscaling via NNPP methodology to yield outputs compromises considerable original map information, resulting in the degraded model performance metrics exhibited in the figures. Yet, as Figure. 5a demonstrates, despite this information loss, the neural network approach can still remarkably shorten the time required to search for the optimal path. Despite sacrificing some fidelity, the neural network method can substantially reduce the computation time for finding the optimal path, as Figure. 5a illustrates.
@@
As shown in Table 2, we tested the model’s path planning performance on a validation set consisting of 1000 maps, wherein ${A^{\star}NN}$ denotes the approach using the procedure of Figure. 6, utilizing the model output to guide ${A^{\star}}$ for path replanning.
@@
Figure. 11 illustrates several scenarios of path planning on a ${256\times 256}$ map. It can be observed that the model’s output effectively avoids hazard terrain (where whiter pixels indicate higher traversal costs, indicating the increased steepness). The model provides relatively safe regions. For certain scenarios, the model presents two alternative paths, indicating its ability to handle semantic information of the global map. During actual planning, the model selects the path with a higher probability value. The corresponding data for the five scenarios are presented in Table 3.
@@
The data presented in the table substantiates the claim that the method proposed in this paper significantly reduces the time consumed in optimal path planning, without substantial compromise on the path’s quality. Specifically, on a ${256\times 256}$ grid map, the NNPP (Neural Network Path Planning) method has achieved a 4.6-fold acceleration in the path searching time of the ${A^{\star}}$ algorithm, with a mere 2.9% decrease in path quality. As the map size increases, the advantages of the NNPP method are expected to become even more pronounced.
5.2 The Effects of Different Gaussian Positional Encoding

### v3

@@ 4.1 Network Structure
In this section, we introduce the neural network structure of the proposed NNPP model. The model takes as input a three-channel combination tensor, where the first channel represents a cost map as detailed in Section 3, with the grayscale value of each pixel indicating the cost of the robot passing through that pixel, with higher costs indicating pixels that are closer to being impassable. The second channel encodes the path planning starting point using a Gaussian representation, as described in Section 4.3. Similarly, the third channel represents the endpoint’s Gaussian encoding. The model’s output is a two-dimensional single-channel probability map, visualized as a color map in Figure LABEL:fig:eps2. Each pixel in the probability map stores a probability value ${p}$ between 0 and 1, indicating the probability that the pixel contains the optimal path. Higher probabilities imply the model is more inclined to believe the optimal path passes through that pixel.
Figure 4: The model architecture diagram. It consists of an encoder, a decoder, and a contextual structure. The input is a three-channel tensor, and the output is a single-channel tensor.
@@
The detailed structure of the model is shown in Figure 4. It uses an encoder-decoder architecture comprising a backbone module, pyramid pooling module and decoder module. From a unique perspective, the role of the model is equivalent to performing semantic segmentation on the map containing the start and end points, segmenting out those pixels that represent feasible paths, while we also expect the model to have a very fast inference speed. Therefore, in designing the model, we drew inspiration from excellent examples of real-time semantic segmentation models.
@@
The NNPP model employs the STDC [26] as the backbone network. Within its multi-stage encoding architecture, the spatial resolution of feature maps is reduced by a factor of two after each STDC stage. As such, a high-level semantic description of the input map can be extracted as the final stage5 output of the STDC backbone. The low-level features output by stage3 and stage4 will be fused with the high-level features in the decoder to enhance the model’s performance for path planning. Feature fusion across different semantic levels has been shown to be effective in improving performance for tasks requiring precise localization such as path planning [18], by incorporating both low-level spatial details and high-level semantic information. Therefore, in our model’s decoder, we fuse the low-level features from earlier stages with the high-level features, in order to leverage spatial details that can help produce more accurate path planning outputs.
@@
To enhance global scene contextualization, PSPNet [27] incorporated a pyramid pooling module (PPM) that concatenates pooling representations at multiple scales prior to the convolution layers, thereby forming both local component features and global contextual representations. Compared to the original PPM, we simplify it to use three different pyramid scales, and remove the concatenate operation between the original feature map and the upsampled features, as in the original PPM. Our simplified PPM can reduce computational cost. Specifically, we reduce the number of pyramid scales from the original 4 to only 3. Previous work has shown that features at different pyramid levels become increasingly semantic [28]. Therefore, a fewer number of pyramid levels, such as 3, can capture sufficient semantic information, while reducing computational cost.
Figure 5: Attention fusing model. The model fuses the low-level features from earlier stages with the high-level features, in order to leverage spatial details that can help produce more accurate path planning outputs.
@@
Feature fusion modules are commonly employed in semantic segmentation networks to enhance feature representations. More specifically, fusing multi-level features is instrumental in achieving high segmentation accuracy. The proposed model draws inspiration from the design of ppliteseg [29], utilizing an Attention Fusion Module to amalgamate low-level features and high-level features from the decoder. The structure of the Attention Fusion Module (AFM) [29] is shown in Figure 5.
@@ 4.2 Dataset Generation
Obtaining the ground truth optimal path on uneven terrain is challenging as it involves semantic segmentation and terrain traversability analysis using intrinsic parameters of the mobile platform. For instance, generating annotated data and training labels for planetary surfaces can be a very time-consuming task [30]. To address this issue, NNPP does not require any manually labeled data. Instead, it uses the optimal path obtained by ${A^{\star}}$ as a label to train the model. The 20-meter resolution lunar DEM dataset from CE2TMap2015 [31] was used to generate a ${256\times 256}$ map. The cost of traversing each pixel was calculated using the method outlined in Eq.1.
The classical ${A^{\star}}$ algorithm is a graph search algorithm commonly utilized for global path planning [32]. It is capable of finding the shortest path for a robot to travel between two predetermined positions in a known static environment. In the classical ${A^{\star}}$ algorithm, as in Eq.5, ${F\left(n\right)}$ represents the cost value of the ${n}$-th node. During the search process of the ${A^{\star}}$ algorithm, the cost values of all neighbor nodes of the current node are calculated, and the node with the smallest value is inserted into the open list.
@@
To facilitate the mobile robot’s automatic avoidance of untraversable areas caused by hazard uneven terrain, this paper defines the cost function of ${A^{\star}}$ to include the traversability cost as follows:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}G\left(n\right)={L_{sn}}+\omega\cdot{T_{sn}}}$ (6)
Where ${{L_{sn}}}$ represents the accumulated Euclidean distance that from the starting grid ${s}$ to the current grid ${n}$, calculated as in Eq.3, and ${T_{sn}}$ represent the accumulated traversability cost from ${s}$ to ${n}$, ${{T_{sn}}=\sum_{k=1}^{n-1}{T_{k,k+1}}}$ and ${T_{k,k+1}}$ is computed as in Eq.2 This paper set ${H\left(n\right)}$ to be the Euclidean heuristic function.
Next, we conducted a comprehensive analysis and established an optimal value for ${\omega}$ in Eq.6. An optimal path should simultaneously satisfy two criteria: (1) minimizing the $CC$ value which is cumulative traverse cost and (2) minimizing the Euclidean path length. When ${\omega=0}$, the algorithm ignores the traversability cost of each grid cell and directly searches for the shortest path from the start point to the end point. In most cases, this shortest path is a straight line. The larger the value of ${\omega}$, the greater the weight of the traversability cost.
To determine the impact of ${\omega}$ on the traversability cost and Euclidean distance cost, we conducted additional experiments as shown in Figure LABEL:fig:r1.
Figure 6: On the same map, with fixed start and end points, the path searching result using different weight values ${\omega}$. For better visualization, the path is shown in bold. The weight values ${\omega}$ from left to right are 0, 0.01, and 0.05.
As the value of ${\omega}$ increases, the path becomes more curved, as shown in the Figure 6. This is to avoid high-cost grid cells and minimize the cumulative traverse cost of the path. Using the cost map in Figure 6 as an example, we fixed the start and end points and varied the ${\omega}$ from 0 to 0.2 to obtain the results shown in Figure 7.
Figure 7: The left figure shows the relationship between ${\omega}$ and the normalized path length ($LEN$) and the normalized accumulated cost ($CC$), while the right figure is a double-y axis line graph. The left axis represents the change of $LEN+CC$ with ${\omega}$ (purple curve), and the right axis represents the time consumed for each path planning (pink curve).
Normalized accumulated Euclidean length and traversability cost of searched paths are denoted by $LEN$ and $CC$, respectively, as shown in the left subfigure of Figure 7. The two values were then added together to obtain the results shown in the right subfigure of Figure 7. As the weight increases from zero, the algorithm returns a larger $LEN$ value and a smaller $CC$ value. The time consumed by the algorithm increases significantly at the beginning and then stabilizes.
There exists a weight value that minimizes the sum of $LEN$ and $CC$ values, which we consider to be the optimal balance between path length and cumulative cost. We selected 100 maps as samples, fixed the start and end points, calculated the optimal ${\omega}$ value for each sample, and obtained the result of ${0.011\pm 0.003}$. When constructing subsequent datasets, we used ${{\omega}=0.011}$ for all samples.
@@ 4.3 Detailed Training Process and Evaluation Metric
Positional encoding is a powerful technique for infusing data with information regarding its position via embedding. In natural language processing (NLP), this approach is often leveraged to ensure that a model is aware of the location of words within a sentence. Inspired by this technique, NNPP makes use of positional encoding to embed location information about the starting and ending points on a map. The resulting encoded output is concatenated with the map, generating a three-channel tensor that serves as the input for the model. After encoding the starting point using sine and cosine functions to represent its absolute position [33], the model’s outputs were not feasible. However, by switching to relative position encoding, more desirable results were achieved. This is because in practical path planning, the system is less concerned with absolute positions and more focused on relative distances between points. Specifically, our interest lies in the relative position of each cell in the elevation map compared to the starting point and the goal point.
@@
In this paper, a two-dimensional Gaussian distribution function centered around the starting point is used as the position encoding, with each grid in the elevation map represented by its corresponding value in the Gaussian function. This value reflects the relative position of the grid with respect to the starting point. The same approach is taken for the ending point. Eq.7 represents the calculation method for the Gaussian distribution function value at ${(x,y)}$. Parameter ${cx}$ represents the ${x}$-coordinate of either the starting or ending point, while ${cy}$ represents the ${y}$-coordinate. The impact of the ${\sigma}$ parameter on model performance was analyzed in the experiments in section 5.2, and the optimal value was determined.
In the field of path planning, scholars have proposed many evaluation methods for path quality [34], among which average path planning time is an important metric that reflects the search efficiency of the algorithm. In elevation maps, the proposed method uses the cumulative traverse cost of a path to measure its optimality, believing that an optimal path should have the minimum cumulative traverse cost.
Figure 8: The segmentation results obtained with different threshold values ${(td)}$. The leftmost image represents the model’s output map, while the remaining images depict the segmentation results obtained using different threshold values. The black regions indicate areas where the pixel values in the model’s output map exceed the ${td}$.
This paper employs a learning-based approach to predict regions within a given map that contain optimal paths, whereas existing evaluation metrics are designed for a complete path rather than a region. Consequently, there is a need for designing an evaluation metric that directly assesses the precision of the model predictions. Figure 8 illustrates the results of segmenting the model outputs according to diverse thresholds ${td}$, with the black areas representing regions assigned probabilities lower than ${td}$ by the model. The designation ${{}^{\prime}connected^{\prime}}$ denotes whether the segmented regions unite the starting point and end point. The present study decreases ${td}$ gradually from high to low values while detecting whether the two points are connected, thereby determining an optimal binary threshold and the corresponding region segmentation map.
Subsequently, the total number of pixels that the model deems to contain optimal paths within the segmented regions is calculated, with the metric evaluated using the following Eq.8. ${Z_{label}}$ represents the area of the labeled path, which is the sequence of grid cells obtained by running the ${A^{\star}}$ algorithm on the traversability cost map, ${Z_{model}}$ denotes the area segmented using the adaptive threshold, and the areas are measured in number of pixels. $ModelMetric$, denoted as $MM$, measures the overlap between the heuristic region output by the NNPP and the ground truth optimal path. A higher $MM$ value indicates better model performance, while a lower value suggests poorer performance.
@@
In summary, there are three evaluation metrics: the average path planning time ${(AT)}$, the cumulative traverse cost of a path (${CC=\sum{{}_{i}^{path}{T_{i}}}}$,${T_{i}}$ represent the traversability cost value of gird ${i}$ in the path grid list, as Eq.1), and ${(MM)}$ as defined in Eq.8.
@@ 5 SIMULATION EXPERIMENTS
In this section, we delve into intricate parameter tuning and hardware environment settings, designing three experiments to demonstrate the efficacy of the proposed method. As detailed in 4.2, a training set of 100,000 samples of resolution ${256\times 256}$ was synthesized, 5000 samples were procured for evaluation, and an additional 1000 samples were sequestered for performance metric validation in the present experiments. The training, elevation and validation sets were mutually exclusive from each other. The optimizer of choice for parameter optimization was AdamW with an initial learning rate of 0.001 and a weight decay of ${{10^{-4}}}$. After training for 150 epochs, the model was able to predict traversable areas between start and end points on 1000 terrain samples that it had not been exposed to previously, showcasing its generalization capabilities.
5.1 Comparative Experiments at Different Scales
The computationally intensive nature of utilizing algorithm ${A^{\star}}$ to search for optimal paths within elevation maps stems from the stipulation that ${A^{\star}}$ algorithm necessitates constructing a global search graph in advance of searching. The purpose of the graph is to store the open-list and closed-list during the search process and maintain the parent node of each node in the list. At each step of the search, the open-list is sorted to select the node with the highest current priority. Once the open-list grows to include the goal node, the parent nodes of the current node can be traced back from the graph to the starting node, thereby returning a path.
Table 1: Comprehensive experimental data on path planning using the ${A^{\star}}$ algorithm under different map resolutions. The path length measured in terms of the total number of pixels traversed along the path.
Map Resolution Graphing Time/s Searching Time/s Graphing Ratio Path Length
$64\times 64$ 0.1854 0.0230 88.96% 57
$128\times 128$ 0.6466 0.1045 86.09% 113
$256\times 256$ 2.7770 0.4715 85.49% 226
$512\times 512$ 11.1785 2.0906 84.24% 452
$1024\times 1024$ 46.4525 8.6014 84.43% 904
@@ 5.1 Comparative Experiments at Different Scales
Figure 9: The left figure illustrates the variation of path search time with map resolution under different methods. The blue color represents the ${A^{\star}}$ search method, the orange color represents the method combining NNPP with ${A^{\star}}$, and the green color represents the model inference time. The model inference time remains nearly constant at 0.058s, indicating that the model can perform inference at a speed of 17 frames per second (FPS). On the right figure, the trend of Model Metric values for the NNPP method is shown as the map resolution changes.
@@
However, the proposed NNPP approach allows the neural network to focus on those regions that provide high value to subsequent path reconstruction algorithms such as ${RRT^{\star}}$ and ${A^{\star}}$. By reducing the spatial redundancy in the search space, the network can learn a more discriminative representation that facilitates more efficient replanning. Taking ${A^{\star}}$ as an example, constructing the search graph only within the heuristic region output by NNPP, rather than a global search graph, as illustrated in Figure 9, can save significant time. Moreover, as the map size increases, the advantages of the neural network approach become more pronounced. By leveraging the ability of the network to focus the search in the most promising areas, the proposed localized search graph yields increasingly superior speedup and scalability compared to exhaustive search-based algorithms as the problem size grows.
[…]
@@
The model was trained on data of size ${256\times 256}$. When evaluating the model’s performance on ${512\times 512}$ maps, as illustrated in Figure 10, the maps were first downsampled via average pooling. Denoted by $s$ and $g$ are the start and goal positions respectively. Subsequently, the method described in 4.3 was applied to obtain the mask of the heuristic region to the model. Finally, these regions were resized back to the original map size via interpolation. This procedure allowed the model trained on smaller maps to generalize well to larger maps, by focusing its search only on the relevant areas while ignoring the rest. The ability to downsample the input during inference and selectively pay attention to important regions enabled by the proposed method was critical for achieving scalability. The search graph was constructed only within the black regions in Figure 10. Subsequently, the ${A^{\star}}$ method was applied on the search graph to compute the start-to-goal path. In comparison, an exhaustive search graph was built on the original map, and the same ${A^{\star}}$ method was utilized to search for the optimal path. As shown in Figure 15, the heuristic region in the first column is notably smaller than the ${A^{\star}}$ algorithm search area in the second column. By constraining the search space to heuristic region, the localized search graph approach achieved significantly higher efficiency and scalability compared to the exhaustive search baseline. The ability to prune irrelevant regions of the map based on the predicted importance masks allowed the search complexity to grow sublinearly with the map size, in contrast to the quadratic complexity of exhaustive search. This demonstrates the effectiveness of the proposed approach in focusing the search in the most promising regions to guide the path planning algorithm.
@@
As shown in the Figure 10, the masked graph making module accepts two inputs: the model output post masking and the original map. In the original map, regions of interest as discerned by the model are retained and delineated in black while the remainder is rendered entirely white, with the white regions signifying impassable obstacles. In this manner, ${A^{\star}}$ search for the optimal path is constrained to a relatively smaller region. Provided that the specified region encompasses the optimum path, ${A^{\star}}$ will return a 100% consistent path matching the given label.
@@
As Figure 11 illustrates, the blue bars represent the computational time incurred by the NNPP method via the workflow in Figure 10 to derive the optimal path whereas the orange bars denote the time consumed by conventional ${A^{\star}}$ search directly on the map. Evidently, as the map size increases, the neural network more decisively shortens the search time. The right sub-figure of Figure 9 indicates the performance metrics measured using Eq.8 and corresponding standard deviations for the NNPP model at diverse resolutions. In larger scenes, employing the upscaling via NNPP methodology to yield outputs compromises considerable original map information, resulting in the degraded $MM$ value. Yet, despite this information loss, the neural network approach can still remarkably shorten the time required to search for the optimal path.
Figure 11: The relationship between the path length obtained by the algorithm and the search time is examined in the validation set. The blue scattered points represent the method combining NNPP with ${A^{\star}}$, while the orange color denotes the traditional ${A^{\star}}$ method. The fitted lines for both methods are also depicted in the figure. It can be observed that as the path length increases, the search time for that path also increases. The blue method exhibits significant superiority over the orange method.
@@
As shown in Table 2, we tested the model’s path planning performance on a validation set consisting of 1000 maps, wherein ${A^{\star}NN}$ denotes the approach using the procedure of Figure 10, utilizing the model output to guide ${A^{\star}}$ for path replanning.
Figure 12: Some examples of path planning on a ${256\times 256}$ map, where the high probability region, indicated by the color cyan, corresponds to the optimal path suggested by the model. The yellow square represents the starting point, while the red pentagram represents the destination.
@@
Figure 12 illustrates several scenarios of path planning on a ${256\times 256}$ map. It can be observed that the model’s output effectively avoids hazard terrain (where whiter pixels indicate higher traversal costs, indicating the increased steepness). The model provides relatively safe regions. For certain scenarios, the model presents two alternative paths, indicating its ability to handle semantic information of the global map. During actual planning, the model selects the path with a higher probability value. The corresponding data for the five scenarios are presented in Table 3.
Table 2: Detailed experimental data is provided for maps with resolutions of ${256\times 256}$ and ${512\times 512}$. In the table, ${AT}$ represents the average path search time, while ${CC}$ represents the cumulative cost of the path in the cost map.
metric ${256\times 256}$ ${512\times 512}$
$A^{\star}$ $A^{\star}NN$ $A^{\star}$ $A^{\star}NN$
$AT$ 2.7234 0.8511 12.1285 2.4103
$CC$ 43.5157 44.8051 89.1457 91.2667
$SuccessRate$ 100% 99% 100% 99.4%
@@
The data presented in the Table 2 substantiates the claim that the method proposed in this paper significantly reduces the time consumed in optimal path planning, without substantial compromise on the path’s quality. Specifically, on a ${256\times 256}$ grid map, the NNPP method has achieved a 3.2-fold acceleration in the path searching time of the ${A^{\star}}$ algorithm, with a mere 2.9% decrease in path quality. As the map size increases, the advantages of the NNPP method are expected to become even more pronounced.
There are many other methods, such as octrees and binary heaps, that can also effectively enhance the efficiency of graphing in ${A^{\star}}$. The NNPP model proposed in this paper does not enhance efficiency by altering the principles of graphing, instead, it narrows the scope of graphing by learning the distribution characteristics of optimal paths in the traversability cost map. Regardless of the graphing method employed, NNPP significantly improves algorithmic search efficiency.
5.2 The Effects of Gaussian Positional Encoding